OpenStudy (anonymous):

Factor the polynomial x^2 - 5xy + 6y^2

OpenStudy (anonymous):

x(x-5)y(1+6y)

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow x^2 - 3xy - 2xy + 6y^2 }\) \(\Large \color{purple}{\rightarrow x(x - 3y) - 2y(x - 3y) }\) \(\Large \color{purple}{\rightarrow (x - 2y)(x - 3y) }\)

OpenStudy (anonymous):

=x^2-4xy+4y^2-xy+2y^2 =(x-2y)^2-y(x-2y) =(x-2y)(x-3y)